OpenStudy (anonymous):

the standard form of a parabla with a vertex (0,0) and directrix at x=5

OpenStudy (anonymous):

 If that's the graph, that means the the p value is 5 and the format is y^2 = -4px y^2 = -4(5)x y^2 = -20x